some boxes of chocolates are 20 dollars each some are 40 dollars and some are 60 dollars each. find the least sum of money with which we can buy an exact number of chocolate of each kind?
1
Expert's answer
2019-09-19T10:36:31-0400

Solution. Find least common multiple for numbers 20, 40, 60. represent numbers through factors


"20=2 \times 2 \times 5"

"40=2 \times 2 \times 2 \times 5"


"60=3 \times 2 \times 2 \times 5"

As result LMC


"LMC=2 \times 2 \times 5 \times 2 \times 3=120 \, dollars"



Answer. 120 dollars.